1) Repeated doubling, in which each doubling occurs in the same amount of time, is a hallmark of linear growth. A) True B) Fa9) A linear function has a constant rate of change and a curved graph. A) False B) 10) The rate of change is equal to the sloFind the slope of the graph and theyy -intercept. Then sketch the graph for values of x between -10 and 10. 15) y= -x+1 A) slDraw a graph of the function and use the graph to answer the question. 16) The cost of tuition at one community college is $317) The general shape of an exponential growth function is a A) rising line B) rising curve C) falling curve D) falling line
Answer #1

1) Repeated doubling, in which each doubling occurs in the same amount of time, is a hallmark of exponential growth

False

Because its geometric growth. In exponential growth, occurs drastically.

2) False

Money in a bank account earning compound interest at an annual percentage rate of 3% is an example of exponential growth.

The growth rate is a relative rate.

3) True

Suppose you had a magic bank account in which your balance doubled each day. Ifyou started with just $1, you’d be a millionaire in less than a month.

. After only twenty days, the balance would be$220= $1,048,576

4)

Linear growth occurs when a quantity grows by the same relative amount in each unit Of time.

5)

Td=4 t=48

2 t/Td = 2(48)/4 = 212